# Large Language Models for Interpretable Mental Health Diagnosis

### v1

@@ 1 Introduction
Recent advancements in large language models (LLMs) suggest their potential in many applications (Friha et al. 2024) including clinical settings (Ullah et al. 2024). Thanks to their excellent processing and understanding of natural language, they can generate diagnostic suggestions based on medical literature and patient data. However, their adoption in clinical settings still faces challenges. For example, LLMs are prone to issues like hallucinations (Huang et al. 2023; Bai et al. 2024), lack of explainability (Zhao et al. 2023), lack of consistency (Moore, Deshpande, and Yang 2024), and limited proficiency in complex reasoning (Huang and Chang 2023). To date, no existing approach effectively combines LLMs with mechanisms that guarantee accuracy and interpretability in the context of mental health diagnosis, thus leaving a critical gap.
To fill the gap, we propose a method that combines LLMs with constraint logic programming (CLP), leading to a practical tool for assisting clinicians in making mental health diagnosis. Specifically, our method leverages LLMs to translate natural language descriptions of mental health diagnostic criteria from manuals such as DSM-5-TR and ICD-11 CDDR to logic rules, thus reducing the cognitive burden on domain experts. Simultaneously, we use an off-the-shelf CLP engine for solving the logic rules to ensure that the diagnostic output is verifiably correct, while enhancing interpretability through the explicitly defined rules and objectives.
Figure 1 shows the overall flow of our method. First, natural language text from a diagnostic manual (e.g., ICD-11 CDDR) is fed into an LLM, which generates a candidate logic program codifying the diagnostic rules. Next, a domain expert manually reviews the code to ensure that the LLM-generated rules accurately encode the manual’s criteria. Finally, the finalized logic program is used by a CLP engine to generate the diagnosis result based on the information of an individual patient. By combining the natural language processing capabilities of the LLM with the logical reasoning capabilities of the CLP engine, our method delivers accurate and inherently interpretable diagnostic outcomes.
@@ 6 Conclusion
We have presented a method that explores the novel application of constraint logic programming for clinical decision support systems (CDSS) in mental health diagnosis and critically evaluates the role of LLMs in this domain. Our evaluation demonstrates that while LLMs show promise in diagnostic tasks, they face significant limitations when used directly, including challenges in consistency, interpretability, and ethical considerations related to patient data. By leveraging LLMs for code generation and combining their capabilities with expert inspection, we propose a safer and more reliable approach that ensures diagnostic logic aligns with clinical criteria. Our findings underscore the importance of systems that balance the efficiency of LLMs with the rigor of expert validation. Future work will explore domain-specific fine-tuning of LLMs, evaluate the approach on real-world datasets, and extend the Datalog encoding to address more nuanced diagnostic criteria and specifiers.

### v2

@@ 1 Introduction
Recent advancements in large language models (LLMs) suggest their potential in many applications (Friha et al. 2024) including clinical settings (Ullah et al. 2024). Thanks to their excellent processing and understanding of natural language, LLMs can generate diagnostic suggestions based on medical literature and patient data. However, their adoption in clinical settings still faces challenges. For example, LLMs are prone to issues like hallucinations (Huang et al. 2023; Bai et al. 2024), lack of explainability (Zhao et al. 2023) and consistency (Moore, Deshpande, and Yang 2024), and limited proficiency in complex reasoning (Huang and Chang 2023). To date, no existing approach effectively combines LLMs with mechanisms that guarantee accuracy and interpretability in the context of mental health diagnosis.
To fill this critical gap, we propose a method that combines LLMs with constraint logic programming (CLP), leading to a practical tool for assisting clinicians in making mental health diagnosis. Specifically, our method leverages LLMs to translate natural language descriptions of mental health diagnostic criteria from manuals such as DSM-5-TR and ICD-11 CDDR to logic rules, thus reducing the cognitive burden on domain experts. Simultaneously, we use an off-the-shelf CLP engine for solving the logic rules to ensure that the diagnostic output is verifiably correct, while enhancing interpretability through the rules and objectives explicitly defined via CLP.
Figure 1 shows the overall flow of our method. First, natural language text from a diagnostic manual (e.g., ICD-11 CDDR) is fed into an LLM, which generates a candidate logic program codifying the diagnostic rules. Next, a domain expert manually reviews the code to ensure that the LLM-generated rules accurately encode the manual’s criteria. Finally, the finalized logic program is used by a CLP engine to generate the diagnosis result based on the information of an individual patient. By combining the natural language processing capabilities of an LLM with the logical reasoning capabilities of a CLP engine, our method delivers accurate and inherently interpretable diagnostic outcomes.
@@ 6 Conclusion
We present a novel approach that integrates large language models (LLMs) with constraint logic programming (CLP) to design a clinical decision support systems (CDSS) for mental health diagnosis. Our evaluation demonstrates that while LLMs show promise for diagnostic tasks, they still face significant limitations when used directly, including issues with consistency, interpretability, and ethical concerns related to patient data. To address these challenges, our method utilizes LLMs to generate logic programs that encode diagnostic rules, and CLP engines to produce diagnostic results based on patient data. We propose that this hybrid approach, combined with expert validation, ensures that diagnostic reasoning is aligned with clinical criteria, enhancing reliability and safety in clinical decision-making. Future work will explore domain-specific fine-tuning of LLMs, evaluate the approach on real-world datasets, and extend the Datalog encoding to address more nuanced diagnostic criteria and specifiers.